Rieger International is attempting to evaluate the feasibility of investing $95,000 in a piece of equipment that has a 5-year life. The firm has estimated the cash inflows associated with the proposalas shown in the following table. The firm has a 12% cost of capital.

Year (t) Cash inflows (CFt)
1 $20,000
2 25,000
3 30,000
4 35,000
5 40,000

a. Calculate the payback period for the proposed investment.
b. Calculate the net present value (NPV) for the proposed investment.
c. Calculate the internal rate of return (IRR), rounded to the nearest whole percent, for the proposed investment.
d. Evaluate the acceptability of the proposed investment using NPV and IRR. What recommendation would you make relative to implementation of the project? Why?
Answer #1
a. Calculate the payback period for the proposed investment.
Payback period : Payback period is the time required for cumulative cash inflows to recover the cash outflows of the project.
Payback period (PBP) = Year before full recovery + (Unrecovered cost at start of year/Cash flow during year)

We see that CFs in 3 yrs = 20000+25000+30000 = 75000
CF is Y4 = 35000. But Cumuulative CFs from Y1 to Y4 = 75000+35000 = 110,000 which is more than Initial Inv. So our PBP is between 3 & 4 yrs.

SO PBP = 3 + (95000-75000)/35000 = 3.57 yrs

b. Calculate the net present value (NPV) for the proposed investment.
NPV = NPV(Rate,CF1,Cf2....CF5) + Initial Inv
ie NPV = NPV(12%,20000,25000,30000,35000,40000) - 95000 = $9,081

c. Calculate the internal rate of return (IRR), rounded to the nearest whole percent, for the proposed investment.
IRR = IRR(CFs) = IRR(-95000,20000,25000,30000,35000,40000) = 15.36%

d. Evaluate the acceptability of the proposed investment using NPV and IRR. What recommendation would you make relative to implementation of the project? Why?
Acceptance criteria for a project is that it should have a positive NPV and IRR should be more than Cost of Capital
Here we have NPV as $9081 & IRR is 15.36% which meets Proj selection criteria.
So we should accept this project